### v1

# AtomGPT: Atomistic Generative Pre-trained Transformer for Forward and Inverse Materials Design

Abstract
Large language models (LLMs) such as generative pretrained transformers (GPTs) have shown potential for various commercial applications, but their applicability for materials design remains underexplored. In this article, we introduce AtomGPT, a model specifically developed for materials design based on transformer architectures, to demonstrate the capability for both atomistic property prediction and structure generation. We show that a combination of chemical and structural text descriptions can efficiently predict material properties with accuracy comparable to graph neural network models, including formation energies, electronic bandgaps from two different methods and superconducting transition temperatures. Furthermore, we demonstrate that AtomGPT can generate atomic structures for tasks such as designing new superconductors, with the predictions validated through density functional theory calculations. This work paves the way for leveraging LLMs in forward and inverse materials design, offering an efficient approach to the discovery and optimization of materials.
keywords:
American Chemical Society,
[…]
IR,NMR,UV
{tocentry}
In recent years, machine learning (ML) has emerged as a powerful tool in materials science, offering the potential to accelerate materials discovery by learning from existing data and making predictions for new materials 1, 2, 3. Such ML models can be used for forward design tasks such as given an atomic structure predict its properties, or inverse design tasks such as given properties predict its atomic structures. Among various ML approaches such as descriptor-based, graph-based and text-based methods, large language models (LLMs) have shown promises to the exploration of vast materials design space beyond the limitations of traditional methods4, 5, 6.
The Generative Pre-trained Transformer (GPT) is a type of large language model (LLM) originally developed for natural language processing and has demonstrated remarkable success in generating coherent and contextually relevant text 7, 8, 9. Models such as ChatGPT 10 have been used for code generation/debugging, literature review and numerous other tasks. However, if we attempt to perform forward/inverse materials design tasks, the outcomes are quite poor. Nevertheless, inspired by its simplicity of usage and the massive success of ChatGPT, we introduce an alternate model, AtomGPT, an atomistic generative pre-trained transformer tailored for forward and inverse materials design.
There have been several previous research works for forward prediction material properties such as descriptor-based methods including models in MatMiner 11, Classical Force-field Inspired Descriptors (CFID) 12; graphs-based methods such as Crystal graph convolutional neural networks (CGCNN) 13, MatErials Graph Network (MEGNet) 14, Atomistic Line Graph Neural Network (ALIGNN) 15; and language-based methods such as LLM-Prop 16 etc. Similarly, some of the earlier inverse design methods are crystal diffusional variational autoencoder (CDVAE) 17, Fourier-transformed Crystal Properties (FTCP) 18, G-SchNet 19, MatBERT 20, Crystal-LLM 21, and xyztransformer 22.
AtomGPT is a deep learning model 6 that leverages a few transformer architectures such as GPT2 23 and quantized Mistral-AI 24 to learn the complex relationships between atomic structures and material properties. We use JARVIS-DFT 25, 26, 27 for training AtomGPT. JARVIS-DFT includes a diverse range of structural, electronic, and thermal properties. By learning from this dataset, AtomGPT acquires a rich representation of the materials space, enabling it to predict properties of unseen materials and generate new material structures with desired properties.
[…]
AtomGPT’s application potential is vast, spanning various domains of materials science. In this work, we focus on designing Bardeen–Cooper–Schrieffer (BCS)-superconductors 28, 29, 30 with relevant material properties for forward and inverse design methods. Finding superconductors with high transition temperatures ($T_{C}$) at ambient as well as extreme conditions has been an active area of research for decades28. Predicting such quantities using experiments and computational methods such as density functional theory 29 can be time-consuming and artificial intelligence-based methods 30, 31 can be helpful to accelerate the design process. In addition to $T_{C}$, some of the other properties of interest for superconductor design are formation energy (should be negative) and electronic bandgap (should be close to zero to follow BCS formalism), which are also considered in this work.
Figure 1: Schematic overview of the AtomGPT workflow. AtomGPT can be used for both forward model (atomic structure to property) predictions as well as inverse design (property to atomic structure generation) using LLMs. a) interconnected text to material property predictions, text input to atomic structure generation, b) an example crystal structure for BCS-superconductor MgB2 (JVASP-1151), c) text description of atomic structure of MgB2 including explicit atomic structure as well as chemical information using ChemNLP32, d) text prompt to explicit atomic structure generation example using Alpaca format 33.
An overview of the AtomGPT package is shown in Fig. 1a. AtomGPT can be used to predict the property of a material given its information or generate its structure given its property information in terms of text. To illustrate these processes, we show an example for MgB2 which is a well-known BCS superconductor with experimental $T_{C}$ of 39 K 34, and DFT prediction of 32.7 K 30 as shown in Fig. 1d. In order to represent the chemical and structural information of materials we use ChemNLP package32, but not limited by it. For example, other packages such as Robocrystallographer 35 could be used here as well. The ChemNLP description includes instructions such as chemical formula, space group, X-ray diffraction peaks, bond-distance, bond-angles and explicit fractional coordinates of a material as shown in Fig. 1c.
First, we discuss the method involved in forward property prediction models. We use the GPT2 model available in the publicly available transformers library 36. AtomGPT allows easy integration of any other related huggingface library models available in the transformers library to be easily used in place of the GPT2 model 23. GPT2 was developed by OpenAI and its successors GPT3.5 and GPT4 are now commercial and have revolutionised the field of artificial intelligence. GPT-2 was primarily trained on CommonCrawl dataset and utilizes a stacked transformer decoder architecture, which consists of multiple layers of transformer blocks 23. Each transformer block contains two main components: a multi-head self-attention mechanism and a position-wise feed-forward network. The input to the model is a sequence of tokens, which are first converted into embeddings and then passed through the transformer blocks. Scaled dot-product attention used in a transformer model can be written as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
where $Q$, $K$, and $V$ represent the query, key, and value matrices, respectively. Here, $d_{k}$ is the dimensionality of the key vectors. The multi-head attention is obtained by concatenating multiple such attention heads. The multi-head self-attention mechanism allows the model to focus on different parts of the input sequence when computing the output for a particular token.
As the LLM models such as GPT2 were primarily designed for text to text models and not explicitly for regression applications, we modify the language model head with a linear neural network model to predict one or several properties as output(s). A pre-trained GPT-2 is loaded as a causal language model (CausalLM), which means it’s designed for tasks where the prediction of each word token depends only on the previous tokens in the sequence (e.g., text generation). We replace the language model head of the GPT-2 model with a new neural network in AtomGPT. The new network is a sequential model consisting of two linear layers. The first linear layer maps the hidden states of the GPT-2 model to a lower-dimensional latent space dimension (here 1024). The second linear layer maps this latent representation to a single output value for a regression task such as formation energy or superconductor $T_{C}$. There are various versions of GPT-2 model, we use GPT2-small with 124 million parameters for forward design tasks. We train such models using V100 GPUs with a batch size of 10, L1 loss function, and a learning rate of 10-3 for 200 epochs. The hyperparameters such as batch size, learning rate, latent space dimension were chosen by varying them for the smallest superconductor database, and used for other tasks with larger datasize such as formation energy and bandgaps.
The dataset used for this work is taken from the JARVIS-DFT and includes formation energies (55713), OptB88vdW bandgaps (55713), Tran Blaha modified Beck-Johnson potential based bandgaps (18167) and superconducting transition temperatures (1058) at ambient pressures 25, 26, 27. The dataset provides atomic structure information such as atomic coordinates, constituent element types and lattice parameters along with DFT calculated properties mentioned above. The dataset is used for 80:10:10 splits for training, validation and testing splits.
As LLM models require text as inputs, we provide a description of an atomic structure using ChemNLP 32 along with explicit lattice parameter and coordinates information which are then fed to the model discussed above. The combined input text of chemical description and atomic information provides a powerful representation of both chemical and structural information of a material.
After training such models, we show the performances Table 1 in the forward design model box for formation energy per atom, OptB88vdW bandgaps, TBmBJ bandgaps and superconducting transition temperatures using corresponding test sets. We compare the model performance for identical train-test splits from CFID, CGCNN and ALIGNN methods. We find that AtomGPT outperforms ALIGNN for OptB88vdW bandgap and $T_{C}$ while for TBmBJ bandgaps and formation energies ALIGNN still dominates other models. The superior performance of ALIGNN for formation energy can be attributed to some of the sophisticated information such as bond-angles, directional graphs and atom feature information 15 which are not explicitly included in the AtomGPT model, Nevertheless, AtomGPT performance for formation energy is closer to CGCNN which is also a GNN model with sophisticated graphs for atomistic systems. Note, it took dedicated efforts to develop such domain specific models such as ALIGNN, while finetuning GPT2 model is very straightforward. For some cases such as TBmBJ, the performances of ALIGNN and AtomGPT are comparable which is promising.
Next, we demonstrate the inverse design task developing a separate model. While the forward prediction model uses GPT2 with a modified large model head, the inverse design model is based on supervised fine-tuning with Mistral-AI 7 billion parameter model 24 using Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning (PEFT) 37. Mistral is a powerful model with 7.3 billion parameters and has been shown to outperform Large Language Model Meta AI (LLaMA) 2 13B 38, LLaMA 1 34B 39, and ChatGPT3.5 10 on several publicly available benchmarks. Some of the key features of Mistal are: sliding window attention to adeptly manage sequences of varying length while minimizing inference cost and grouped-query attention for swift inference (decrease inference time), details of which can be found in Ref. 24.
Forward models
Prop/MAE CFID CGCNN ALIGNN AtomGPT
$E_{form}$ (eV/atom) 0.142 0.063 0.033 0.072
OPT $E_{g}$ (eV) 0.299 0.199 0.142 0.139
MBJ $E_{g}$ (eV) 0.531 0.407 0.310 0.319
$T_{C}$ (K) 1.99 2.60 2.03 1.34
Inverse models
Database/RMSE CDVAE AtomGPT
SuperConDB 0.24 0.08
Carbon24 0.36 0.32
Table 1: Performance measurement for forward and inverse design tasks and comparison with Classical Force-field Inspired Descriptors (CFID)12, Crystal Graph Convolutional Neural Network (CGCNN)13, About Atomistic Line Graph Neural Network (ALIGNN)15 and Crystal Diffusion Variational Autoencoder (CDVAE)17 models. Best performing model results are shown as bold texts. The forward property prediction (Prop.) models are shown using mean absolute error (MAE) metric between held out test set and property predictions, while the inverse models are compared using reconstruction loss in root mean square distance between held out test structures and those using the generative models.
Figure 2: Chemical prototype and density of (a–b) the AtomGPT generated structures and (c–d) the 1058 structures used for training from the JARVIS-DFT database.
We specifically focus on the superconductor inverse design task. This requires finetuning LLM models such as Mistral AI. Such finetuning requires transforming the instructions into a specialized protocol such as Alpaca 33. An example of converting MgB2 into alpaca is shown in Fig. 1d. It consists of a python dictionary with Instruction, Input and Output keys. As directly finetuning such a model can be computationally expensive, we use Parameter-Efficient Fine-Tuning (PEFT) method using Hugging Face ecosystem (transformers, PEFT, Transformer Reinforcement Learning (TRL)). We patch the Mistral model with fast LoRA (Locally Reweighted Attention) 37 weights for reduced memory training. After obtaining the PEFT model, corresponding tokenizer, and alpaca dataset (with 80 % training, 10 % validation and 10 % test splits for the superconductor dataset), we instantiate a supervised fine-tuning with a batch size of 10, AdamW 8 bit optimizer and cross-entropy loss function. This loss function measures the difference between the predicted probability distribution over the vocabulary and the true distribution (i.e., the one-hot encoded target words). After the model is trained, we evaluate the model on the test set with respect to reconstruction/test performance. To further clarify, after training the model on the train set, we keep the instruction and input keys in the test set and let the trained model generate outputs. We parse the outputs to create corresponding crystal structures. Then we use the StructureMatcher algorithm 40, which finds the best match between two structures considering all invariances of materials. The root mean square squared error (RMSE) is averaged overall matched materials. Because the inter-atomic distances can vary significantly for different materials, the RMSE is normalized following the works in Ref. 17. Note that this is just one of the metrics for generative models for atomic structures and there can be numerous other types of metrics. It is important to calculate the $T_{C}$ using a high fidelity method such as DFT later that acts as a more reliable metric for the generation process. We show the performance of the model for the JARVIS-SuperconDB (1058) and Carbon24 (10153) datasets 41. We notice from Table 1 that AtomGPT has lower RMSE than CDVAE for both SuperconDB and Carbon24 databases suggesting that given a specific prompt, the model can generate structures very similar to the held out test set. Note that only one property $T_{C}$ and specific type of instructions are used here, and LLMs allow easy addition of other properties and conditional statements as well.
Figure 3: Density functional theory calculation results for one the identified structures. a) atomic structure visualization, b) Eliashberg function for Nb2B structure with a predicted $T_{C}$ of 4.76 K.
Next, we generate materials using such models for identifying potential candidate superconductors. We focus on Borides (but are not limited by it) as they are one of the systems of interest for our future experimental synthesis. One of the first computational discoveries and experimentally synthesized superconductors is FeB4 42 which further motivates such application. We attempt to generate systems with the formula XaBb, XaYbBc, XaYbZcBd, where X,Y,Z are periodic table elements other than Boron and a,b,c can be integers such as 1,2,3. After generating such 890 structures, we visualize the chemical prototypes and density distributions in Fig. 2a and Fig. 2b. We observe that most of the generated structures belong to prototypes A2B, A, AB and AB2. Their densities can attain values upto 15 gcm-3. The prototypes generated are similar to those existing in the training dataset i.e., the JARVIS-DFT superconductor dataset as shown in Fig. 2c. The densities of the generated structures also closely resemble distributions of the training set with a peak around 5 gcm-3. We use the formation energy property model developed above to check if they have negative formation energy, bandgap models to make sure they have close to zero bandgaps and $T_{C}$ greater than 1 K leading to 30 candidate materials. Then, we subject these candidates for ALIGNN-FF (developed in Ref. 43) based pre-optimization. After the pre-optimization, we subject these materials to JARVIS-DFT workflow using Vienna Ab initio Simulation Package 44, 45 and OptB88vdW functional. We now perform electron-phonon calculations for these optimized structures. Note that this step is computationally expensive and can take days if not weeks for each material. One of the potential candidates we further investigated is Nb2B. We find the optimized lattice lengths as 3.19 Å, 3.18 Å and 4.52 Å and lattice angles as 93.7°, 90°and 90°with monoclinic spacegroup of P2/m. It has a density of 7.1 gcm-3 which is almost 3 times higher than Silicon (2.25 gcm-3) and aluminum (2.66 gcm-3). We calculate the formation energy and energy above the convex hull of this system as -0.13 eV/atom and 0.3 eV respectively. Finally, we perform McMillan-Allen-Dynes formula based superconductor transition temperature calculations using quantum espresso using workflow developed in Ref. 30 to calculate its $T_{C}$.
We follow the workflow used to generate the training data (in ref. 30), where we performed electron phonon calculations (EPC) calculations using non-spin polarized DFT-perturbation theory (PT) 46, 47 (using the interpolated/Gaussian broadening method 48) with the Quantum Espresso (QE) software package 49, PBEsol functional 50, and the GBRV 51 pseudopotentials. The EPC parameter is derived from spectral function ${\alpha}^{2}F(\omega)$ which is calculated as follows:
${\alpha}^{2}F(\omega)=\frac{1}{2{\pi}N({\epsilon_{F}})}\sum_{qj}\frac{\gamma_{qj}}{\omega_{qj}}\delta(\omega-\omega_{qj})w(q)$ (2)
where $\omega_{qj}$ is the mode frequency, $N({\epsilon_{F}})$ is the density of states (DOS) at the Fermi level ${\epsilon_{F}}$, $\delta$ is the Dirac-delta function, $w(q)$ is the weight of the $q$ point, $\gamma_{qj}$ is the linewidth of a phonon mode $j$ at wave vector $q$ and is given by:
$\gamma_{qj}=2\pi\omega_{qj}\sum_{nm}\int\frac{d^{3}k}{\Omega_{BZ}}|g_{kn,k+qm}^{j}|^{2}\delta(\epsilon_{kn}-\epsilon_{F})\delta(\epsilon_{k+qm}-\epsilon_{F})$ (3)
Here, the integral is over the first Brillouin zone, $\epsilon_{kn}$ and $\epsilon_{k+qm}$ are the DFT eigenvalues with wavevector $k$ and $k+q$ within the $n$th and $m$th bands respectively, $g_{kn,k+qm}^{j}$ is the electron-phonon matrix element. $\gamma_{qj}$ is related to the mode EPC parameter $\lambda_{qj}$ by:
$\lambda_{qj}=\frac{\gamma_{qj}}{\pi hN(\epsilon_{F})\omega_{qj}^{2}}$ (4)
Now, the EPC parameter is given by:
[…]
with $w(q)$ as the weight of a $q$ point. The superconducting transition temperature, $T_{c}$ can then be approximated using McMillan-Allen-Dynes 52 equation as follows:
$T_{c}=\frac{\omega_{log}}{1.2}\exp\bigg{[}-\frac{1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\bigg{]}$ (6)
where
[…]
In Eq. 6, the parameter $\mu^{*}$ is the effective Coulomb potential parameter, which we take as 0.1. It is important to note that the robustness of this workflow was heavily benchmarked against experimental data and higher levels of theory in ref. 30, 53, 54, 55, which indicates that the training data for these deep learning models is of high quality, given the level of theory used to produce the data.
We find a $T_{C}$ of 4.76 K for Nb2B and observe no negative modes in the phonon density of states curve suggesting dynamical stability. Similar DFT calculations can be performed for other structures as we populate more candidates, which can then be added to the JARVIS-DFT Superconductor database 30, 53, 54, 55. Moreover, we can extend this workflow to include other candidate systems, such as nitrides and carbides. As the database of superconductors expands, both forward and inverse models have the potential for further improvement, which is necessary for enhanced predictions. After DFT validation, we can recommend these computationally discovered candidates for experimental synthesis. Superconductor design remains one of the most fascinating areas of research in physics and chemistry. Our work here serves as a proof-of-concept that demonstrates the potential of LLM models, such as GPTs, to accelerate superconductor materials design. Moreover, it can be easily extended to other classes of materials such as alloys, semiconductors, themoelectrics, dielectrics, piezoelectrics, solar cells etc.
Acknowledgements
We thank computational resources from the National Institute of Standards and Technology (NIST). This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Please note commercial software is identified to specify procedures. Such identification does not imply recommendation by NIST.